OpenStudy (anonymous):

A=[ 7 -5; 5 -9] B= [ -8 4; 6 -8] What is AB??

OpenStudy (anonymous):

\[A=\left( \begin{array}{cc} 7 & -5 \\ 5 & -9 \end{array} \right),B=\left( \begin{array}{cc} -8 & 4 \\ 6 & -8 \end{array} \right)\]

OpenStudy (anonymous):

First row * First column 7*-8 + -5*6 =-86

OpenStudy (anonymous):

\[\left( \begin{array}{cc} -86 & \square \\ \square & \square \end{array} \right)\]

OpenStudy (anonymous):

Now for second one. first row * second colum 7*4 +-5*-8=68 \[\left( \begin{array}{cc} -86 & 68 \\ \square & \square \end{array} \right)\]

OpenStudy (anonymous):

can you do the last two?

OpenStudy (anonymous):

yes i think so, thank you!